Circular No. 8185

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

PECULIAR VARIABLE IN CRUX H. E. Bond, Space Telescope Science Institute, reports that spectra

(resolution 0.22 nm; range 402–493 nm) of Tabur’s possible nova (cf. IAUC 8184), obtained with the SMARTS Consortium 1.5-m telescope at Cerro Tololo by S. Gonzalez on Aug. 21.95 UT, appear to confirm that the object is a classical nova. Although taken under poor observing conditions, the spectra show weak Hβ emission with a broad, blue-shifted absorption component, characteristic of a nova before or near maximum light.

M. Della Valle, Osservatorio Astrofisico di Arcetri, Florence; D. Hutsemekers, University of Liege; and I. Saviane and E. Wenderoth, European Southern Observatory (ESO), write: “Preliminary analysis of a spectrum (range 380–740 nm; resolution ∼ 0.1 nm) of the variable star reported on IAUC 8184, obtained on Aug. 22.02 UT with the ESO 3.6-m telescope (+ EFOSC2) at La Silla, shows a reddened continuum characterized by a lack of the typical signatures exhibited by Fe- or He/N-type classical novae in their early evolution. We detect faint Hα and Hβ emission lines flanked by P-Cyg profiles. The expansion velocity derived from the minima of the P-Cyg profiles is ∼ 500 km/s. All of this suggests this object not to be a classical nova, but possibly a post-asymptotic-giant-branch star in flare-up, a V4332 Sgr-type object (Martini et al. 1999, A.J. 118, 1034), or a V838 Mon-type object (IAUC 7785; Bond et al. 2003, Nature 422, 405) caught at an early stage.”

Visual magnitude estimates: Aug. 21.944 UT, 9.0: (W. Souza, São Paulo, Brazil); 21.948, 9.1 (J. G. S. Aguiar, Campinas, S.P., Brazil); 22.493, 10.2 (A. Pearce, Nedlands, W. Australia).

SUPERNOVA 2003hm IN UGC 2295 Further to IAUC 8184, T. Boles reports the discovery of an apparent

supernova (mag 17.7) on unfiltered CCD images taken on Aug. 20.112 and 21.069 UT with a 0.35-m reflector in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 2^{h}48^{m}58^{s}.39, δ = +3^{o}10′07′′.6, which is ≈ 6′′.4 west of the center of UGC 2295. SN 2003hm is not present on Boles’ images from Jan. 7 and Feb. 17 (limiting mag 19.0), and it is not present on Palomar Sky Survey 1990 red or 1992 blue plates.

2003 August 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green